OpenStudy (anonymous):

Solve the equation to find a positive value of c: 62 + 82 = c2

OpenStudy (anonymous):

c=72 All you have to do for this equation is add it then divide by two 62+82=144/2=72 :D

OpenStudy (anonymous):

thanks! I was being lazy I guess...

OpenStudy (anonymous):

It's fine! Sometimes things look harder then they are :D Good luck!!

OpenStudy (anonymous):

Oh, wait, I forgot, when I copied and pasted, there was an error. The equation is ctually 6 to the second + 8 to the second = c to the second do you know how to sovle this?

OpenStudy (anonymous):

Uhh I can try!! Give me a second!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

14 is my guess. You might wanna have some one validate or invalidate my answer. Sorry!

OpenStudy (anonymous):

Ok, thanks anyway! :)